Table 2 shows a list of 10 generators which have submitted offers for supplying electricity and their prices Table 2. List of generators and their details GeneratorMin Generation Max GenerationPrice (E/MWh) Number 2 4 300 1000 300 800 250 200 500 320 500 200 1200 20.5 100 20 17 5 50 200 80 100 10 25 7 27 15 12 10 (i) Arrange the generators in merit order Determine which generators would be called upon to supply with the predicted demand given in Table 3 (ii) (ii) List the marginal prices which would be used in each two hour period (iv) Calculate the daily revenue of each generator Table 3. Predicted demand for each two-hour period Two hour period Predicted Demand 0:01-2:00 2:01-4:00 4:01-6:00 6:01-8:00 8:01-10:00 10:01-12:00 12:01-14:00 14:01-16:00 16:01-18:00 18:01-20:00 20:01-22:00 22:01-24:00 2000 1800 2000 3800 4000 4000 3600 3400 4500 5200 4200 3000
Answer #1

Answers (i): The Generators are promising to supply various levels of electricity at variable prices. On the basis of performance, i.e. providing more of the electric supply at cheapest prices with most variable option is the basis of the below selection of the merit list of the Generators:

1又 2 j O 叉OD ,250 1 5 5 50 10 1 070

Answers (ii): When there is demand of electricity, the most preferable option will be to choose the generator which could provide with the maximum supply of electricity at the minimum cost per Mega What per hour. Therefore, as the below table shows, that Generator 1 will be preferred through out most of the times in the day when the demand for electricity is on the higher side. Generators 2 will be demanded when the demand is comparatively lower, and Generator 3 will be demanded when the demand for electricity will be the lowest in the day:

wo how 1xa 1 800 1132 3960 41000 4000 X-2 120 12.01 一 15; 14 0-16: 6 16,01 18:01 知:6P 3400 45 0b 35211 1113 1-01 이 의이ㅡ 01

Answer (iii) : The marginal price of the Generators have been shown above on 2 hourly basis, with Generator 1 having been utilized during the peak hours when its total income on a two hourly basis would be (12 x 2 = 24 ). Similarly the overall marginal prices of the generator are shown in the above figure 2.

Answer (iv): The Daily revenue of each Generator would be as follows :

Generator 1 = 5 x ( 12*2 ) = 120

  Generator 2= 2 x ( 9*2 ) = 36

  Generator 3 = 5 x ( 11*2 ) = 110

The rest of the generators being not beneficial will not be called upon and therefore will not make any profit.